Calculate the pH of the solution made by adding 0.50 mol of HObr in 0.30 mol of KOBr to 1.00 L of water. The value of Ka for HObr is 2.0 * 10^-9
Answer #1
Concepts and reason

The concept used to solve this problem is based on the chemical equilibrium.

The pH of solution is used to determine acidity and basicity of a solution.

Fundamentals

The pH of solution can be determine by Henderson-Hasselbalch equation as follow.

pH=pKa+log[conjugatebase][acid]pH = p{K_{\rm{a}}} + \log \frac{{\left[ {{\rm{conjugate base}}} \right]}}{{\left[ {{\rm{acid}}} \right]}}

Here pKa{\rm{p}}{{\rm{K}}_{\rm{a}}} is written as follow.

pKa=logKap{K_{\rm{a}}} = - \log {K_{\rm{a}}}

Here, Ka{K_{\rm{a}}} is acid dissociation constant.

The pKap{K_{\rm{a}}} of the solution can be calculated by formula as follow.

pKa=logKap{K_{\rm{a}}} = - \log {K_{\rm{a}}}

Substitute 2.0×1092.0 \times {10^{ - 9}} for Ka{K_{\rm{a}}} .

pKa=log(2.0×109)=8.7\begin{array}{c}\\p{K_{\rm{a}}} = - \log \left( {2.0 \times {{10}^{ - 9}}} \right)\\\\ = 8.7\\\end{array}

Here, weak acid is HOBr{\rm{HOBr}} and its conjugate base is OBr{\rm{OB}}{{\rm{r}}^ - } . In the buffer mixture, HOBr{\rm{HOBr}} is act as conjugate base.

The Henderson-Hasselbalch equation is as follow.

pH=pKa+log[conjugatebase][acid]pH = p{K_{\rm{a}}} + \log \frac{{\left[ {{\rm{conjugate base}}} \right]}}{{\left[ {{\rm{acid}}} \right]}}

Thus, Substitute 8.7 for pKap{K_{\rm{a}}} , 0.30molL10.30{\rm{ mol }}{{\rm{L}}^{ - 1}} for [conjugatebase]\left[ {{\rm{conjugate base}}} \right] and 0.50molL10.50{\rm{ mol }}{{\rm{L}}^{ - 1}} for [weakacid]\left[ {{\rm{weak acid}}} \right] .

pH=8.7+log(0.300.50)=8.5\begin{array}{c}\\pH = 8.7 + \log \left( {\frac{{{\rm{0}}{\rm{.30}}}}{{{\rm{0}}{\rm{.50}}}}} \right)\\\\ = 8.5\\\end{array}

Ans:

The pH value of buffer solution is 8.5.